Given that f(x)=6x+11 and g(x)=x^3, find;(f*g)(-2)
Answer by Alan3354(69443)

Given that f(x)=6x+11 and g(x)=x^3, find;(f*g)(-2)
---------
(f*g)(x) = (6x+11)*x^3 = 6x^4 + 11x^3
(f*g)(-2) = 6(-2)^4 + 11(-2)^3 = 96 - 88 = 8
------------------
Or,
f(-2) = -1
g(-2) = -8
(f*g)(-2) = 8